OpenStudy (anonymous):

a merchant has coffee with$50 a pound that she wishes to mix with 90 pounds of coffee with $80 a pound to get a mixture that can be sold for $60 a pound . how many pounds of the $50 coffee should be used?

OpenStudy (anonymous):

(50)x+90(80)=(60)(x+80)

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

Imran has made a slight error in his reply, it shud b as follows: Let x pound of $50 be mixed with 90 pound of $ 80 so x(50) + 90 (80) = (x + 90) (60) 50x + 7200 = 60x + 5400 7200 - 5400 = 60x - 50x 1800 = 10x 1800/10 = x so x = 180 Hence he shud mix 180 pounds of $50 coffee

OpenStudy (anonymous):

If yr question has been answered to yr satisfaction, pls click on "Good Answer" button to close the question